OpenStudy (abbycross167):

I will FAN AND GIVE A MEDAL! Can someone please help me redo this problem? Please? I really need help... Below I'll add the question then my answer then what my teacher said.

OpenStudy (abbycross167):

Question~ Write an equation that models the different amounts (in pounds) of hamburger and chicken you can buy. 9. You are buying hamburger and chicken for a barbecue. The hamburger costs $3.52 per pound and the chicken costs $4.57 per pound. You have $40 to spend. My Answer~ Hamburger~ Round $3.52= $3.50 $3.50*5.5lbs = $19.25 Chicken~ 4.50*4.5lbs= $20.25 Cost of hamburger added to the cost chicken~ $19.25+$20.25= $39.50 I can buy 4 ½ lbs of chicken and 5 ½ lbs of hamburger.

OpenStudy (abbycross167):

What My Teacher said~ Where is the equation that models this situation? Look at the example below... Example: You are in charge of buying fruit cups and yogurt for the school picture. A box of fruit cups costs $4.00 and a box of yogurt containers costs $3.00. You have $48 to spend. Write an equation that models the different combinations of boxes of fruit cups and yogurt containers you can buy. Solution: The key concept here is to remember: Value times Quantity. Fruit cups: 4.00x Yogurt containers: 3.00y Adding those together gives the total to spend: 4x + 3y = 48.

OpenStudy (jessicaloraine):

Ok. I see what you did. Why did you round $3.52 at the beginning?

OpenStudy (abbycross167):

umm I thought that was what I needed to do

OpenStudy (jessicaloraine):

so your teacher wrote 4x+3y=48 because 4 times the number of fruit cups and 3 times the number of yogurt containers is 48

OpenStudy (jessicaloraine):

ok but you dont need to round anything at all. in fact, you do not need to calculate anything

OpenStudy (abbycross167):

oh ok

OpenStudy (jessicaloraine):

so the hamburger costs $3.52 and the chiken cost $4.57 you do not need to think about the per pound that is written in there you actually only need to look at the main words not hamburger or chiken so 3.52 times x numbers of hamburger plus 4.57 times the number of chicken =40 3.52x+4.57x=40

OpenStudy (jessicaloraine):

no need to round anything or any form of calculation

OpenStudy (abbycross167):

oooh I see now! So for the answer to this question all I have to put is--- Hamburger: 3.52x Chicken: 4.57x Adding those together gives the total to spend: 3.52x+4.57x=40 Does that look right? @JeruDavis